OpenStudy (anonymous):

2/3=x/6=9/y To find the value of y in this proportion you must divide 54 by A)2 B)3 C)4 D)6

OpenStudy (aravindg):

from equality 2*6=3*x

OpenStudy (aravindg):

12=3x x=?

OpenStudy (aravindg):

similarly 2/3=9/y so do as above